OpenStudy (anonymous):

The area of a triangle is 24 square inches. What is the height of the triangle if the base length is 6 inches? 4 inches 8 inches 18 inches 20 inches

OpenStudy (anonymous):

@DaBest21

jimthompson5910 (jim_thompson5910):

Area of triangle = (base*height)/2 A = (b*h)/2

jimthompson5910 (jim_thompson5910):

A = 24 b = 6 h = unknown A = (b*h)/2 24 = (6*h)/2 24*2 = 6*h ... Multiply both sides by 2. 6*h = 48 h = ???

OpenStudy (anonymous):

8

jimthompson5910 (jim_thompson5910):

yes h = 8

OpenStudy (anonymous):

dabest21 you can help me on my next one i putting up right now

OpenStudy (anonymous):

8 bc A=b*h/2 so 6*8=48 48/2=24 @liana1026

OpenStudy (anonymous):

sure